OpenStudy (anonymous):

please help I'm having trouble figuring out finding the inverse of a function on this problemf(x)=5x+3over 4

OpenStudy (anonymous):

f(x) is y so put the equation equal to y then make x the subject when you do interchange x and y and that should be the inverse

OpenStudy (anonymous):

let me know what you got

OpenStudy (anonymous):

so x= 5y+3over 4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

jus to be sure the original equation is f(x) = 5x + 3/4 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right so replace f(x) with y so the equation will be y= 5x +3/4

OpenStudy (anonymous):

y - 3/4 = 5x

OpenStudy (anonymous):

(y - 3/4) all over 5 = x x is now the subject . now where you see x you put y and where you see y you put x so the inverse of f(x) = x- 3/4 all over 5

OpenStudy (anonymous):

ok i understand now thank you

OpenStudy (anonymous):

your welcome